### v1

# AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents

Abstract
AI agents have drawn increasing attention mostly on their ability to perceive environments, understand tasks, and autonomously achieve goals. To advance research on AI agents in mobile scenarios, we introduce the Android Multi-annotation EXpo (AMEX), a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents. Their capabilities of completing complex tasks by directly interacting with the graphical user interface (GUI) on mobile devices are trained and evaluated with the proposed dataset. AMEX comprises over 104K high-resolution screenshots from 110 popular mobile applications, which are annotated at multiple levels. Unlike existing mobile device-control datasets, e.g., MoTIF [2], AitW [19], etc., AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions, each averaging 13 steps with stepwise GUI-action chains. We develop this dataset from a more instructive and detailed perspective, complementing the general settings of existing datasets. Additionally, we develop a baseline model SPHINX Agent and compare its performance across state-of-the-art agents trained on other datasets. To facilitate further research, we open-source our dataset, models, and relevant evaluation tools. The project is available at https://yuxiangchai.github.io/AMEX/.
1 Introduction
AI assistants on mobile devices, such as Siri on iPhones, Bixby on Samsung devices, and Xiao AI on Xiaomi smartphones, have become increasingly prevalent in recent years. While these assistants are adept at managing various routine tasks like setting alarms, conducting web searches, and reporting weather conditions, their capabilities are predominantly confined to interacting with system-built applications. Furthermore, although many can interface with third-party apps via APIs, the lack of universal API support across different mobile operating systems often hampers their performance.
In contrast, human users can complete tasks on mobile devices purely based on visual information from the screen. Inspired by this, researchers are exploring alternative approaches [3, 10, 30] that process vision and natural language inputs, specifically screenshots in mobile environments. We refer to these as Mobile GUI-Control Agents, or GUI Agents for short. These agents are designed to manipulate the user interface elements on the screen directly, with the abilities to interpret natural language commands and analyze screenshot layouts and element functionalities, which theoretically enable agents to execute any task on any app.
Figure 1: An example of a screenshot-instruction (Blue tab) pair illustrating the multi-level annotation of AMEX. Red boxes + Brown tabs: selected GUI interactive elements and their corresponding functionalities. Green tab: the detailed annotation of the element. Yellow tab: the description of the entire screenshot. Purple hand icon + tab: the current action and the annotation.
Table 1: Comparison of AMEX to other datasets. Scale: the number of unique instructions on general third-party apps, average steps per instruction and screenshots. Diversity: screenshot description, element labels, element functionality and the action details for stepwise operation.
Dataset # Unique
General Inst. # Avg
Steps # Screen-
shots Screen
Desc. Screen
Element Element
Func. Action
Detail
PixelHelp 187 4.2 $\sim$800
UGIF 480 6.3 $\sim$3.3K
MoTIF 480 4.5 $\sim$21K
AitW 1539 6.5 $\sim$510K *
AitZ 2504 7.5 $\sim$18K * $\blacksquare$
AMEX 2946 12.8 $\sim$104K
  • * indicates elements are mis-annotated and the bounding boxes are not well-aligned.
  • $\blacksquare$ indicates containing only action results of the element interacted at each step.
However, the latest GUI agents, such as CogAgent [10] and AppAgent [28], exhibit unsatisfactory performance when handling real-world tasks. The primary issues with these agents are their lack of awareness of page layouts and their limited understanding of the functionalities of various user interface elements. These shortcomings are largely due to the absence of a comprehensive, detailed, large-scale dataset. Recent research has introduced several related datasets for the Android operating system, featuring different types of annotations. For instance, AitW [19] is a large-scale dataset annotated with instructional operations and screenshot icon detection. However, the data quality, especially in element-level annotation which is constructed by a pretrained IconNet, is subpar due to the unregulated annotation style. Besides, only 2.8% of instructions are for general third-party apps, not representative of general usage. Additionally, AitW suffers from heavy data redundancy as 94.9% instructions, including both multi-step and single-step, are in the category of “Web-shopping”. To address these issues, AitZ [32] selects a smaller subset of AitW and re-annotates it with screenshot descriptions and Chain-of-Action-Thought (CoAT) annotations. Though the CoAT paradigm is of some help, the dataset scale shrinks substantially to total 18K screen-action annotation pairs. Other datasets [2, 14, 23] also have various limitations, including the number and practicability of instructions, diversity, accuracy and reliability of annotations, coverage of apps and Android versions. Furthermore, these datasets primarily offer supplemental tree-based representations, such as the View Hierarchy. Notably, the View Hierarchy is not universally available across all apps, nor is it readily accessible to average mobile users. This limitation restricts the practical usability and generalizability of the data, as it does not reflect the typical interaction scenarios encountered by most users.
When human users receive an instruction, they perform several steps beyond merely interpreting the instruction. A human user would comprehend the overall page layout and analyze the current page to identify interactive elements, determining which elements can be tapped and where to scroll. The functionality of those elements are then assessed, particularly the clickable ones. For example, a bell icon likely indicates navigation to a notification page, while a magnifier icon might signify a search page or activate a search bar. Equipped with this information and contextual knowledge, the user can effectively break down the instruction into specific actions for multiple screens.
To align an GUI agent with humans, we design a comprehensive multi-level annotation dataset for the Android OS, named the Android Multi-annotation EXpo (AMEX). AMEX consists of three levels of annotations (see Figure 1): (i) GUI interactive element grounding, (ii) GUI screen and element functionality descriptions, and (iii) instructions with GUI-action chains. In total, AMEX consists of annotations on over 104K high-resolution screenshots, 711K element-wise functionalities, and around 3,000 unique instructions with stepwise GUI-action chains, with an average of 13 steps per instruction (see Table 1 and Table 2). The bounding boxes of GUI interactive elements are filtered and verified by human annotators. Screen descriptions and element functionalities, generated by GPT with built-in app descriptions, are also manually checked. Given random apps and instructions, annotators, all trained with specific and precise guidelines for each action, conduct the operations in a natural manner to achieve the objective. The data is collected from 110 third-party apps with Android 13 on Pixel 7 Pro and Samsung S10 emulators, ensuring up-to-date and prevalent screen resolutions and aspect ratios. The key features of AMEX can be summarized as providing knowledge of the modern mobile GUI environment from multiple levels, and complex human logics and operations on third-party apps.
Our contributions can be summarized as follows: (a) We collect and release AMEX for training and evaluating generalist mobile GUI agents, which has multi-level annotations, providing reliable understandings and complex instructions of the smartphone UI environment; (b) We release the SOTA agent SPHINX Agent, which can serve as the baseline model for future researches on GUI agents.
2 Related Work
GUI-Control Datasets. Table 1 compares several popular GUI-control datasets on Android. While some works [5, 15, 20] focus on the web platform, on Android OS, many works [19, 29, 33] have focused on identifying various types of GUI elements, often assigning numerous classes to different elements. Other studies [2, 14, 23] primarily emphasize action-observation pairs during instructional operations, but their annotations are limited and often require supplemental View Hierarchy (VH) data for each screenshot. Furthermore, these instructions are typically too simplistic for real-world tasks. AitW [19] provides both screen GUI element annotations and instructions, but it includes only a small portion of instructions for third-party apps, with most operations conducted on Chrome and other system-built apps. Additionally, each instruction is repeated multiple times, resulting in significant data redundancy. It also relies on the pre-trained IconNet for automatic annotations, which unfortunately leads to numerous mis-annotated types and misaligned bounding boxes. In response, AitZ [32] filters AitW thoroughly, selecting 2.5K unique instructions and episodes, and introduces the Chain-of-Action-Thought framework to annotate action results and page descriptions better. Despite this refinement, AitZ contains only 18K screen-action pairs.
GUI-Control Agents. Recent advancements have leveraged the extensive world knowledge and robust embodied capabilities of Large Language Models (LLMs) [1, 8, 9, 11, 12, 22] for complex task planning and reasoning within GUIs. A notable approach involves employing business-level generalist models like GPT-4v directly as GUI-control agents. Works [28, 34] have utilized these models, employing extensive prompt engineering to guide the LLM in executing complex tasks. However, the effectiveness of these methods is inherently limited by the capabilities of the available generalist MLLMs [4, 13, 16, 24] and requires meticulous prompt design to achieve optimal results. Alternatively, another research line focuses on fine-tuning smaller LLMs on GUI-specific datasets to imbue them with domain-specific knowledge, thereby enhancing their operational efficiency. For example, CogAgent [10] enhances performance in GUI-related tasks by integrating a high-resolution cross-module that fuses image features from various levels. Similarly, MobileAgent [6] improves sample handling and structuring of input data to make it more consistent and LLM-friendly. Auto-UI [30] utilizes a multimodal encoder-decoder language model based on T5 and BLIP and CoCo-Agent [17] takes element OCR layouts as additional input to enhance the performance. Expanding beyond traditional mobile phone GUI domains, recent studies [26, 31] are exploring the potential of LLM-Agents in operating system-level task execution.
3 Android Multi-annotation EXpo (AMEX)
(a) Element annotation on AitW (left) and AMEX (right).
(b) Demo of GUI interactive elements.
Figure 2: Demonstrations of element annotations of AitW and AMEX. (a) Element bounding boxes in AitW (left) and AMEX (right). Boxes in AMEX are well aligned but boxes in AitW might be misaligned and mis-annotated. (b) GUI interactive elements in AMEX. Red boxes: clickable elements. Blue boxes: horizontally scrollable elements. Yellow box: vertically scrollable element.
Figure 3: Overview of the data collection pipeline. The raw data is from two subsets collected by human annotators and an autonomous tool respectively, while annotators record the GUI-action chains simultaneously. Then raw data is sent to annotators to filter GUI element bounding boxes, then they with raw screenshots are sent as input to GPT to extract the GUI screen and element descriptions, which are then manually checked by humans.
When receiving an instruction, a human user first analyzes the overall screen layout to form a basic understanding of the current Android environment. The user then identifies interactive elements and areas, and assesses the functionalities of those elements. Finally, the user breaks down the instruction into simple, step-by-step actions on each screen. Based on this human cognitive process, we design three levels of annotations in our proposed AMEX training set: (i) GUI interactive element grounding (see Section 3.2), (ii) GUI screen and element descriptions (see Section 3.3), and (iii) instructions with GUI-action chains (see Section 3.4).
The raw data is collected using Android emulators, specifically Android Virtual Device and Genymotion [^1] https://www.genymotion.com/. We developed tools to autonomously perform emulator operations as well as record human actions. These tools leverage Appium [^2] https://appium.io/docs/en/latest/, an open-source, cross-platform test automation tool for Android, iOS, and web apps. Appium collects device screenshots during operations, each of which also corresponds to an Extensible Markup Language (XML) file recording the basic screen layout with element attributes, such as bounding boxes and in-app descriptions.
3.1 Data Collection Pipeline
An overview of the data collection pipeline is illustrated in Figure 3. The raw data is collected through two methods: human instruction-following GUI manipulations and autonomous GUI controls. Human GUI manipulations involve recording stepwise operations for each instruction, and simultaneously storing screenshots and each screen’s XML data before each stepwise operation. In parallel, an autonomous script controls emulators to collect additional screenshots and their XMLs. These two subsets comprise the entire raw dataset. Then for each screenshot, initial bounding boxes of interactive elements and their in-app descriptions (if available) are parsed from the corresponding XML. Human annotators then review each screenshot to filter out all the misaligned boxes, which serve as the interactive element grounding annotations. With the in-app descriptions, GPT generates the functionalities of the selected elements and provides descriptions for the whole screenshot. Human annotators then further check the descriptions of functionalities. More detailed methods are discussed in the following sections and Appendix A.1.3.
3.2 Level I: GUI Interactive Element Grounding
Existing datasets [19, 32] typically classify elements on the screen, such as icons, texts, and images, based on their types. Instead of adhering to the traditional classification paradigm, we define interactive elements more broadly as any elements that users can interact with, regardless of their specific types (see Figure 2(a)). Specifically, interactive elements in our dataset are only categorized into two subsets: (i) clickable elements and (ii) scrollable elements.
Clickable Elements are the most common components in a screen. They typically include clickable icons, images, texts, and compounds that combine several categories. Figure 2(b) illustrates various clickable elements in red boxes. We also include certain “typeable” elements, such as search bars, because most typeable elements require a prior click action to enable typing.
Scrollable Elements typically occupy larger areas on the screen. In most cases, a scrollable element area supports a pair of actions, such as “scroll down” and “scroll up” or “scroll left” and “scroll right.” Generally, the vertical scrollable area forms the main frame in the middle of the screen, while the horizontal scrollable area serves as a “carousel” for listing different categories or activities. Figure 2(b) illustrates these two types of scrollable elements.
As mentioned in Section 3.1, the collection of screenshots and their XML data utilizes two methods. Here, we provide a more detailed pipeline for executing an autonomous script to perform operations on an Android emulator. The script is designed to traverse an app in an unconstrained manner and collect data. It performs actions (see Appendix A.1.1) at regular time intervals to allow each page to fully load. After waiting, the script captures the current screenshot along with the corresponding XML file. To ensure efficiency and prevent redundancy, we limit the maximum number of operation steps per script execution. Additionally, a break mechanism is implemented to exit potential dead loops, ensuring the script progresses effectively. Upon the raw data (yellow part in Figure 3), we extract bounding boxes for interactive elements from the XML associated with each screenshot. While scrollable elements are typically parsed with high quality, clickable elements are more problematic. This issue is from the fact that app developers often design elements with overlapping or layered coverage, leading to the generation of bounding boxes even for elements that are visually hidden from the user’s view. To address this, annotators manually inspect each screenshot, identifying and selecting only the interactive elements that are actually visible within the interface (see Appendix A.1.2).
3.3 Level II: GUI Screen and Element Functionality Descriptions
Previous works on GUI elements often rely on predefined class names to convey the underlying meaning of each element. However, this classification-based method has significant limitations. For instance, a small triangle symbol typically indicates “play” in video or music interfaces, but it might also appear as part of a content creator’s profile image. In such cases, the model erroneously interprets the symbol and repeatedly attempts to tap the icon to play the video. This issue arises because the class-based annotation approach focuses on class labels rather than truly understanding the functionality of each element in the surrounding context.
To ensure the dataset is truly instructive rather than merely icon detection, we focus on describing screen status and element functionalities. Consider the above mentioned example: instead of providing a bounding box for the small triangle and labeling it as “ICON_PLAY,” we present the actual functionality of the element within its context (e.g., “Opens the creator’s personal page”). This strategy offers clear, instructive, and detailed description, enhancing the dataset’s applicability. We adopt GPT-4o for the element-wise annotation to describe screens and interpret element functionalities. Given the compact and dense nature of GUI elements, we apply Set of Markers (SoM) techniques [27] to bolster GPT-4o’s capability for visual localization. The coordinates of these elements are obtained and cleaned in the preceding phases (Section 3.2). Additionally, recognizing that some elements possess abstract functionalities that are challenging to discern, we supplement the prompts for GPT-4o with in-app descriptions (if available) extracted from XMLs to enhance GPT-4o’s comprehension of the screen. After the data generation, human annotators manually check 20% of the descriptions of the element functionality. Further elaboration on this methodology is provided in Appendix A.1.3.
3.4 Level III: Instructions with GUI-Action Chains
Our focus for GUI agents is predominantly on third-party applications rather than system-built apps. We select an average of three apps per category from the Google Play Store. Due to the significant increase in app functionalities over the past years, we estimate that a sample size of approximately 50 apps can adequately represent most common use scenarios.
The instruction generation process comprises three phases. Initially, human annotators create 5-10 complex instructions for each target app, considering its specific purposes and capabilities. These initial instructions, combined with relevant app metadata collected online, serve as input for ChatGPT. It then generates a larger set of 80-100 instructions that exhibit similar structure and intent to the human-provided examples. For example, if a human annotator provides the instruction “Open Google Maps. Find the shortest route from the current location to Empire State Building by car drive,” ChatGPT might produce “Open Google Maps. Find the fastest route from the current location to Rockefeller Center by public transit.” However, due to ChatGPT’s limitations on understanding real-world constraints, human filtering is adopted to remove any unreasonable or impractical instructions. Notably, instructions within AMEX are more complex than those in the AitW dataset, with an average step count of 12.8, nearly double that of AitW.
We define our action space for stepwise GUI operations similarly to AitW: {TAP, SCROLL, TYPE, PRESS_BACK, PRESS_HOME, PRESS_ENTER, TASK_COMPLETE, TASK_IMPOSSIBLE}. TAP actions are characterized by identical touch and lift coordinates, while SCROLL actions involve distinct touch and lift coordinates. TYPE actions are annotated with a type_text attribute specifying the input text. The three PRESS actions correspond to system-level button presses (back, home, enter). TASK_COMPLETE and TASK_IMPOSSIBLE serve as terminal flags for instructions. Additionally, for instructions that involve information query (e.g., “What is the lowest price of the men’s belt?”), we associate the TASK_COMPLETE action with a region of interest, which is defined as the bounding box of the area on the current screenshot, where the answer is expected to appear (see examples in Appendix A.3). This comprehensive action space allows us to fully simulate a wide range of typical use cases.
The collection tool is designed to simultaneously record the screen and action pair together with the XML file. The details of GUI-action chain collection can be found in Appendix A.1.4.
3.5 AMEX Statistics and Test Set
Table 2: AMEX statistics
# Screenshots # Interactive
Elements # Functionalities # Instructions # Avg. Steps # Apps
103,684 1,658,439 711,823 2,946 12.8 110
Figure 4: (a) The number of screenshots by the number of GUI clickable elements. (b) The number of instructions by the number of GUI operations steps. (c) The word cloud of the instructions.
Data statistics. Table 2 lists the statistics of the AMEX dataset. AMEX has approximately 104K screenshots, 1.6M GUI interactive elements for grounding, 712K GUI element functionality descriptions, and 3K instructions. Figure 4 illustrates the distribution of GUI interactive elements per screen and GUI operation steps along with the word cloud of the instructions.
Test set. Instructions and their annotated screen-action chains from 9 out of 55 apps are chosen as the test set, including those on Booking, Gmail, NBC News (News), SHEIN, Citymapper (CM), Microsoft To Do (ToDo), Signal, Yelp and YouTube Music (Music), covering different scenarios. The number of test instructions is 362, around 10% of the total instructions.
The evaluation of AMEX test set is similar to the existing dataset AITW [19]. We perform the action matching to get the partial action match score. The partial action matching score is the number of correct actions divided by the episode length [14]. To ensure the correctness of the test set, we carefully review the trajectory annotations in the unseen test set, where we drop the former step of PRESS_BACK, which indicates the wrong operation of the annotator. However, we keep the PRESS_BACK step to evaluate the agent’s self-correction ability.
4 SPHINX-GUI Agent
Integrating data from AMEX with the existing AitW dataset, we train the SPHINX-GUI Agent tailored for GUI-related tasks as the baseline for AMEX dataset. This Agent is initialized from SPHINX [7] and is specifically trained to generate actions to manipulate the GUI effectively, which are different from the more flexible output formats typical of standard MLLM tasks such as QA. To incorporate GUI-specific intelligence into SPHINX, we have implemented several key modifications:
Instruction Dataset Construction. We transform the AMEX data and the publicly available AitW dataset into a VQA format conducive to instruction understanding, thereby adapting the MLLM as a GUI agent. To embed GUI-specific knowledge within the MLLM, we developed four distinct VQA tasks, detailed in the instruction templates provided in Table 3.
Table 3: Task template examples.
Tasks Examples of Task Templates Source
Screen Description
User: Provide a one-sentence caption for the provided GUI image.
Agent: This page is a film section of the IMDb app, displaying a list of movies or shows with sorting and filtering options.
AMEX
Element Grounding
User: Identify all clickable elements and provide their 2D bounding boxes.
Agent: List of Elements with BBoxes.
AMEX
Functionality Description
User: What is the function of the icon at the coordinates: BBox.
Agent: Click to expand or collapse the sort options.
AMEX
GUI-Action Prediction
User: Please execute the task described by the following instruction: Delete the first email in the inbox.
Agent: CLICK[[0.59, 0.92]].
AMEX & AitW
Structured Representation. To effectively guide the agent in task completion, we incorporate the goal into the input prompt. To prevent redundant actions, we also include a history of actions formatted as $X_{\text{history}}=\left[a_{t-k},\ldots,a_{t}\right]$, where each $a$ denotes an action type, accompanied by a touch point for actions such as SCROLL or TAP. The parameter $k$ denotes the length of the history. Action trajectories from both AitW and AMEX datasets are converted into this uniform representation.
Given the extensive nature of GUI screenshots and the small relative size of some target elements within the screen, we adopt the “any resolution” approach from SPHINX [7]. This method initially partitions the input image into sub-images, then encodes separately. The LLM processes all the resulting visual tokens to ensure comprehensive understanding and interaction capabilities. We follow all other default training receipts in [7] to obtain the final agent.
5 Experiments
5.1 Evaluation Setup and Compared LVLMs
We evaluate different GUI agents including SPHINX Agent developed in Section 4 on AitW and our AMEX dataset respectively. For AitW, we simultaneously train SPHINX Agent on all the subsets and then assess all test sets. We downsample GoogleApps subset to $10\%$ to avoid data imbalance. The training and test split setting follows the one in [19]. For AMEX, we evaluated our SPHINX agents on the test set as described in Section 3.5. Our models are benchmarked against several agents including Auto-UI [30], SeeClick [3], CogAgent [10] and CoCo-Agent [17].
5.2 Experiment Results
The experiment outcomes for AitW and AMEX are presented in Tables 4 and 5 respectively. As indicated in Table 4, SPHINX Agent (SphAgent) trained on AitW exhibit competitive performance on the AitW dataset relative to the baselines. Notably, the introduction of the AMEX dataset enhanced the overall performance by approximately 2.5%, with significant gains observed in the “General” (5%), “Single” (5%) and “WebShopping” (2%) tasks, which indicates the strong complement from AMEX dataset. The GUI element functionality descriptions are served as the “Single” step instructions which strongly promote the performance in the “Single” tasks, while 3K instructions from general third-party apps also boost the performance in the “General” tasks. “WebShopping” is the category with the most complicated tasks in AitW, which also benefits from the complex instructions provided by AMEX.
Table 4: Experiment results on AitW. + indicates CoCo-Agent takes screen element OCR layouts as additional inputs.
Agent # Params Training Data General Install G-Apps Single WebShopping Overall
Auto-UI 5B AitW 68.24 76.89 71.37 84.58 70.26 74.27
[…]
CoCo-Agent+ 7B AitW 70.96 81.46 76.45 91.41 75.00 79.05
SphAgent 7B AitW 68.2 80.5 73.3 85.4 74 76.28
SphAgent 7B AitW + AMEX 73.1 80.5 73.4 90.8 75.8 78.72
Table 5: Experiment results on AMEX
Agent Training Data Gmail Booking Music SHEIN News CM ToDo Signal Yelp Overall
[…]
SphAgent AitW 32.1 45.9 46.1 35.1 48.3 61.1 55.9 43.3 42.9 45.63
SphAgent AMEX 61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.71
SphAgent AitW + AMEX 62.4 68.1 76.3 71.9 68.6 67.3 77.6 66.0 64.1 69.14
5.3 Cross-Domain Experiments
We also examine the generalization capabilities of our agent across different domains. Specifically, the agent trained exclusively on the AitW dataset is tested on the AMEX dataset. The results, as detailed in the first row of Table 5, reveal significant findings. Agent trained only on AitW, shows satisfactory performance on the AitW dataset itself. However, when evaluated on the AMEX dataset, there is a notable decrease in performance. This performance degradation can likely be attributed to several factors: 1) Higher complexity in AMEX. The AMEX dataset generally involves longer operational sequences compared to those in AitW, presenting a higher complexity level that the agent might not have been exposed to during training. 2) Domain gap. There are substantial differences in the visual textures and language instructions between the two datasets. These variations could hinder the agent’s ability to effectively transfer its learned knowledge to a new, unfamiliar domain. Additionally, we observed a significant performance degradation for the SeeClick agent, originally trained on AitW. Beyond the previously mentioned reasons, this decline may also stem from the fact that the original SeeClick model was designed to process only $224\times 224$ image inputs. In contrast, the AMEX dataset comprises high-resolution images and requires interactions with tiny icons. The high-definition images in AMEX introduce new challenges for GUI agents.
5.4 AitW Human Evaluation Results
Table 6: Human evaluation results on 5%-subsampled AitW test sub-set. “MMR” stands for “Mis-Match Random” subset.
Subset Action Sources General Install G-Apps Single WebShopping Overall
MMR SphAgent 90.91 83.65 87.50 87.68 82.39 86.43
AitW Anno 88.18 75.00 82.69 84.78 75.57 81.25
Random SphAgent 92.31 93.75 94.33 94.83 94.55 93.95
AitW Anno 94.23 90.34 93.62 93.97 95.00 93.43
As specified in AitZ [32], several types of error cases are identified in the AitW test set (see Figure5). To assess the unreliability of the original AitW test set and its evaluation methods, we propose AitW-HE (Human Evaluation), a refined subset of the AitW test set. Human annotators evaluate two subsets derived from the original test set. One subset is randomly chosen from episodes where SphAgent receives low scores (Mis-Match Random), i.e., about 66 score compared to the overall 78.72 shown in Table 4, indicating a high mismatch between inference results and AitW annotations. The other subset is randomly selected from the remaining episodes (Random). Annotators first filter out repeated and redundant screenshots (see Figure 5(a)). For each remaining screenshot, they then record whether the SphAgent-inferred action and the original AitW annotation are correct. Figure 6 illustrates cases where annotators mark both the inferred action and the original annotation as correct, even though they interact with different elements.
(a) Repeating useless screenshots and actions.
(b) Wrong task complete.
Figure 5: Error cases in AitW test set. Red dots indicate the actions from the AitW annotations.
Table 6 presents the accuracy scores from human evaluation. The table indicates that the low SphAgent scores in the MMR subset are primarily due to unsatisfactory annotations (AitW Anno), which are used as ground truth during the evaluation. Both SphAgent and AitW Anno in the “MMR” subset have lower scores than those in the “Random” subset, highlighting that the tasks in the MMR subset are relatively more challenging. Furthermore, SphAgent achieves better human evaluation results than the original annotations, demonstrating its effectiveness and close alignment with human judgment. Comparing the results from Table 4 and Table 6, the notable differences in overall scores (i.e., 78.72%, 86.43%, and 93.95%) underscore the unreliability and misleading nature of the AitW test set and its evaluation methods.
Figure 6: Purple dots indicate the SphAgent inferred actions and red dots indicate the AitW annotations. Human annotators mark them both correct even though they are clicking different elements.
5.5 Ablation Study on Multi-level Annotations
To validate the effectiveness of the multi-level annotations proposed in the AMEX dataset, we conducted an ablation study. As shown in Table 7, the agent achieved a performance gain of approximately 1.0 when aided by L1 annotations. When trained with L1 and L3 annotations, the agent exhibited a performance gain of about 1.6 compared to the baseline. The gain from L2 annotations is approximately 2.3. Furthermore, training the agent with the complete set of annotations resulted in a performance gain of approximately 3.5. The ablation study results demonstrate that each level of annotations in AMEX enhances the final agent’s performance.
Table 7: Ablation study of each level of annotations.
Tasks Gmail Booking Music SHEIN NBC CM ToDo Signal Yelp Avg
L1 L2 L3
45.9 64.5 74.4 71.8 70.3 67.4 79.3 64.9 66.3 67.2
58.6 68.3 74.4 68.8 68.1 61.1 79.3 73.2 67.4 68.8
60.3 66.1 77.2 70.2 72.0 64.0 78.1 72.3 65.7 69.5
61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.7
6 Discussions
6.1 Limitations and Future Work
Multi-lingual    Most existing datasets are limited to English, with UGIF [23] being a notable exception, as it includes instructions and screenshots in eight languages. The AMEX dataset contains a small number of screenshots in Chinese and Spanish, primarily due to strict registration and login requirements for Chinese apps and a lack of expertise in other languages. Future work should incorporate multi-lingual screenshots, functionalities, and instructions to create a more robust and comprehensive multi-lingual environment for GUI agents.
Evaluation    The current evaluation method employed in both AMEX and AitW is straightforward: the model predicts the current action based on given instructions, previous operations, and screenshots. However, this evaluation approach has significant limitations in real-world scenarios. It fails to account for loading pages, waiting times for page loading, etc. An ideal evaluation would be an “Arena” (a mobile version of the environment in WebArena [35]) — an independent local virtual device equipped with Android and static background apps. This setup would ensure consistent results for the same queries, avoiding variations due to differing dates, prices, and other dynamic data.
6.2 Ethical Considerations
  • The accounts registered and logged in are all for testing purposes, not including any personal information. The dataset doesn’t contain any private or personal information.
  • The dataset, if misused, could be exploited for undesirable purposes, such as anti-fraud mechanisms and anti-script verification codes (see Appendix A.4), potentially leading to harm.
  • Annotators received remuneration in line with local wage standards for their annotation works.
7 Conclusion
As AI agents become more prevalent, mobile GUI agents are emerging as a research hotspot. To address the lack of fundamental understanding of GUI elements in existing datasets, we present the Android Multi-annotation EXpo (AMEX) dataset, which includes three levels of annotation to provide a more instructive and detailed understanding of UI screens and elements. Additionally, we introduce a state-of-the-art SPHINX Agent, which can serve as a baseline model for future research.
Appendix A Appendix / supplemental material
A.1 Pipeline Details
A.1.1 Autonomous script details
The autonomous script controls the emulator using three actions: TAP, SCROLL, TYPE.
  • For the TAP action, we employ two algorithms. The first randomly selects a clickable element on the current screen, while the second computes the index of a clickable element using a formula to ensure the elements chosen are likely unique. We apply one of these algorithms randomly for different executions.
  • For the SCROLL action, we classify whether an area is vertically or horizontally scrollable by setting a width-height ratio threshold, $\mathcal{R}$. If the element’s ratio exceeds $\mathcal{R}$, it is considered horizontally scrollable; otherwise, it is vertically scrollable. We then randomly select a scrollable element on the current screen and perform a scroll action based on its type.
  • For the TYPE action, we pre-define a list of phrases relevant to the category of apps being tested. For example, [“Women’s dress”, “Nike sneakers”, …] for clothing shopping apps. These phrases are primarily used in search scenarios.
A.1.2 GUI clickable element filtering
Figure 7: Demonstration of human annotator filtering. (a) before filtering. (b) after filtering.
The raw XML information sometimes contains the elements that are covered by other elements or layers. Figure 7 illustrates the same image before and after filtering. Blue boxes in Figure 7(a) are those elements under the current active layer and they still show up from XML parsing. Thus we need human annotators to filter out these blocked elements to get Figure 7(b).
A.1.3 GUI screen and element functionality description collection details
We apply the Set-of-Mark (SoM) [27] technique when using GPT-4o as the description generator. The SoM technique is a visual prompting method designed to enhance the visual grounding capabilities of large multimodal models (LMMs), such as GPT-4V, by overlaying visual marks on image regions. This involves partitioning an image into semantically meaningful regions and adding distinct marks (e.g., alphanumeric characters, masks, or boxes) to these regions. It demonstrates significant improvements in precision and accuracy over traditional prompting methods and other state-of-the-art models. Figure 8 shows the screenshot with SoM technique.
Figure 8: Demonstrations of SoM technique on screenshots.
The prompt is in the following format.
Based on the screenshot of an Android mobile phone from the APP_NAME, please follow the instructions:
1. Understand the Page Content:
- Analyze the overall content of the page.
- Provide a brief summary of the page content in 1-2 sentences.
2. Explain Highlighted Areas:
- Each highlighted area is either clickable or scrollable.
- Treat each highlighted area as a unique and separate entity, using identifiers such as <Region 1> etc.
- If the highlighted area is a general icon, provide its type first in the format ICON_Magnifying_Glass.
- If the highlighted area is more complex, provide a brief description in the format Element(’a poster of
the movie named <La La Land>’).
- Explain the purpose or functionality of each highlighted area. In other words, what result will happen
or what’s the user’s intention when the marked <@ area is clicked or scrolled?
- Some functionality may require an overall analysis, and try to give the functionality specifically and
related to the current screenshot.
3. Additional Information for Highlighted Areas:
There are total NUMBER elements to annotate.
MarkerInformation
4. Output Format:
The output should be in JSON format as follows:
"overall_page_content": "1-2 sentences summarizing the page content",
"Region 1": "ICON_XXX_XXX <functionality>: xxxxx",
"Region 2": "Element(’a poster of a movie xxxx’)
<functionality>: click to see the details and forward the purchase page of the movie xx",
Listing 1: An example prompt for guiding GPT4o to generate the element functionalities for the given Screenshot.
A.1.4 GUI-Action chain collection details
Human annotators are each assigned a random selection of apps and their associated instructions, which they are asked to complete in a natural manner. In contrast to the AitW dataset, our collection methodology allows annotators to make errors and take incorrect steps, leading to a greater prevalence of PRESS_BACK actions. This is motivated by our observation that agents trained on existing datasets exhibit difficulty navigating back to previous pages due to insufficient experience with the PRESS_BACK action. Additionally, after completing an information query task, annotators are asked to manually mark the region of interest on the screenshot using our annotation tool.
A.1.5 Collection resources details
  • GUI interactive element grounding takes approximately 3000 human-hour to filter bounding boxes described in Appendix A.1.2.
  • GUI screen and element descriptions use GPT-4o API, which consumes about 600 dollars.
  • Instructions with GUI-action chains take approximately 200 human-hour.
A.2 Experiment details
In our implementation, we utilize the internlm-7b variant of the SPHINX-X model, as detailed in [7]. The pre-trained checkpoint for this model was sourced from the official repository mentioned in [21]. For image processing, the input images, each sized $1024\times 1024$, are segmented into sub-images. Visual features from these sub-images are extracted using two distinct visual encoders: DINOv2 [18] and ConvNext [25]. To ensure compatibility in feature dimensions across different modules, linear projection layers are employed to align the channel dimensions. Regarding the model’s parameter settings, as outlined in Section 4, we configure the history window size to four. Additionally, we introduce a special token, <ICON>, specifically designed to identify interactive elements within the interface, strengthening the model’s interpretability and responsiveness to user interactions. The agent is trained on a cluster with 3 nodes, each with eight NVIDIA A100 (80GB) GPUs. The fine-tuning was completed in four epochs.
A.3 More AMEX examples
See Figure 9 for more examples of GUI interactive elements grounding and description.
See Figure 10 and Figure 11 for more examples of instruction with GUI-action chains.
A.4 Examples of Ethical Problems
Figure 12 shows examples of anti-script mechanism where the agent can correctly enter the verification codes.
Figure 9: More examples
Figure 10: More examples
Figure 11: More examples
Figure 12: Demonstration of anti-script mechanism where agents can enter the verification codes correctly.

### v2

@@
 AI agents have drawn increasing attention mostly on their ability to perceive environments, understand tasks, and autonomously achieve goals.
-To advance research on AI agents in mobile scenarios, we introduce the Android Multi-annotation EXpo (AMEX), a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents.
-Their capabilities of completing complex tasks by directly interacting with the graphical user interface (GUI) on mobile devices are trained and evaluated with the proposed dataset.
-AMEX comprises over 104K high-resolution screenshots from 110 popular mobile applications, which are annotated at multiple levels.
-Unlike existing mobile device-control datasets, e.g., MoTIF [2], AitW [19], etc., AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions, each averaging 13 steps with stepwise GUI-action chains.
+To advance research on AI agents in mobile scenarios, we introduce the Android Multi-annotation EXpo (AMEX), a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents which are capable of completing tasks by directly interacting with the graphical user interface (GUI) on mobile devices.
+AMEX comprises over 104K high-resolution screenshots from popular mobile applications, which are annotated at multiple levels.
+Unlike existing GUI-related datasets, e.g., Rico Deka et al.
+(2017), AitW Rawles et al.
+(2024b), etc., AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions with stepwise GUI-action chains.
 We develop this dataset from a more instructive and detailed perspective, complementing the general settings of existing datasets.
-Additionally, we develop a baseline model SPHINX Agent and compare its performance across state-of-the-art agents trained on other datasets.
-To facilitate further research, we open-source our dataset, models, and relevant evaluation tools.
-The project is available at https://yuxiangchai.github.io/AMEX/.
+Additionally, we finetune a baseline model SPHINX Agent and illustrate the effectiveness of AMEX.
+AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents
+Yuxiang Chai1*, Siyuan Huang2*, Yazhe Niu1, Han Xiao1, Liang Liu3, Guozhi Wang3, Dingyu Zhang1, Shuai Ren 3, Hongsheng Li 1, 1MMLab @ CUHK, 2SJTU, 3vivo AI Lab, *Equal contribution, Corresponding author: hsli@ee.cuhk.edu.hk
 1 Introduction
-AI assistants on mobile devices, such as Siri on iPhones, Bixby on Samsung devices, and Xiao AI on Xiaomi smartphones, have become increasingly prevalent in recent years.
-While these assistants are adept at managing various routine tasks like setting alarms, conducting web searches, and reporting weather conditions, their capabilities are predominantly confined to interacting with system-built applications.
-Furthermore, although many can interface with third-party apps via APIs, the lack of universal API support across different mobile operating systems often hampers their performance.
-In contrast, human users can complete tasks on mobile devices purely based on visual information from the screen.
-Inspired by this, researchers are exploring alternative approaches [3, 10, 30] that process vision and natural language inputs, specifically screenshots in mobile environments.
-We refer to these as Mobile GUI-Control Agents, or GUI Agents for short.
-These agents are designed to manipulate the user interface elements on the screen directly, with the abilities to interpret natural language commands and analyze screenshot layouts and element functionalities, which theoretically enable agents to execute any task on any app.
+In recent years, AI-powered virtual assistants, such as Siri, Bixby, and Xiao AI, have evolved as key tools for facilitating interactions between users and mobile devices.
+These assistants have demonstrated to be effective in managing routine tasks such as setting alarms, performing web searches, and reporting weather conditions.
+However, their functionality is often restricted to system-built applications or third-party apps supported by application programming interfaces (APIs).
+This limitation highlights a significant gap between the capabilities of current AI assistants and the diverse, open-ended ways humans interact with mobile devices.
+Unlike AI assistants, human users can accomplish tasks on mobile devices by relying solely on visual information from the screen.
+Humans intuitively interpret graphical user interfaces (GUIs), analyze page layouts, and infer the functionalities of interactive elements to complete tasks in various apps.
+Inspired by this human ability, researchers are exploring new paradigms for mobile interaction that leverage vision and natural language processing.
+One such paradigm involves the development of Mobile GUI-Control Agents, or GUI Agents, which aim to directly interact with screen elements based on visual and textual inputs, such as screenshots and natural language instructions.
+These agents hold the potential to transcend the limitations of traditional API-based assistants, enabling more flexible and universal interactions with any app or mobile interface.
+Existing GUI agents show promise but face significant challenges in real-world applications.
+Their effectiveness is possibly limited by a lack of understanding of GUI layouts and the functionalities of interactive elements through the observation of wrong predictions.
+These limitations largely arise from the absence of comprehensive datasets that adequately reflect the complexity and diversity of mobile GUI environments.
+While several instructional datasets Burns et al.
+(2021); Sunkara et al.
+(2022); Gubbi Venkatesh et al.
+(2024); Rawles et al.
+(2024b) have been introduced to address this issue, they struggle with limitations such as inaccurate annotations, insufficient task diversity, and a lack of representativeness for general mobile usage.
+Similarly, other efforts Deka et al.
+(2017); Li et al.
+(2020); Feng et al.
+(2024); Bunian et al.
+(2021) that focus on GUI element annotation are constrained by limited dataset scale, outdated app versions, and traditional annotation styles that do not provide sufficient information for GUI agents.
 Figure 1: An example of a screenshot-instruction (Blue tab) pair illustrating the multi-level annotation of AMEX.
@@
 Purple hand icon + tab: the current action and the annotation.
-Table 1: Comparison of AMEX to other datasets.
-Scale: the number of unique instructions on general third-party apps, average steps per instruction and screenshots.
-Diversity: screenshot description, element labels, element functionality and the action details for stepwise operation.
-Dataset # Unique
-General Inst. # Avg
-Steps # Screen-
-shots Screen
+Human users approach GUI-based tasks by integrating multiple cognitive processes.
+They interpret the overall layout of the page, identify interactive elements and assess their functionality, such as recognizing that a bell icon likely leads to notifications and a magnifier icon likely triggers a search function.
+Users then decompose natural language instructions into actionable steps, navigating across multiple screens to achieve their goals.
+To design effective GUI agents, it is essential to replicate these human cognitive processes, which requires datasets that comprehensively capture the structure, functionality, and interaction patterns of mobile GUIs.
+To address this need, we introduce a new dataset, the Android Multi-annotation EXpo (AMEX), specifically designed to advance the development of GUI agents by providing a multi-level understanding of mobile GUIs.
+AMEX includes three levels of annotations: (i) GUI interactive element grounding, (ii) GUI screen and element functionality descriptions, and (iii) instructions with GUI-action chains.
+The dataset comprises over 104K high-resolution screenshots, 21K screen descriptions with 300K element-wise functionalities, and approximately 3,000 unique complex instructions, with an average of 12.8 steps (see Figure 1, Table 1, and Table 2).
+To ensure annotation precision and quality, all annotations, including bounding boxes, screen and element functionalities, and instructions with GUI-action chains, are verified by human annotators trained with detailed guidelines.
+Our contributions can be summarized as follows: (a) We collect and release AMEX, a multi-level dataset, providing reliable and detailed understandings of the smartphone GUI environment, which can serve as a strong supplementary dataset to boost the performance of GUI agents; (b) We train SPHINX Agents, which can serve as the baseline models for future researches on GUI agents and illustrates the effectiveness of AMEX.
+2 Related Work
+2.1 GUI-related Datasets
+Dataset Screen
 Desc. Screen
 Element Element
-Func. Action
-Detail
-PixelHelp 187 4.2 $\sim$800
-UGIF 480 6.3 $\sim$3.3K
-MoTIF 480 4.5 $\sim$21K
-AitW 1539 6.5 $\sim$510K *
-AitZ 2504 7.5 $\sim$18K * $\blacksquare$
-AMEX 2946 12.8 $\sim$104K
+Func. Task &
+Action # Screen-
+shots # Element
+Func. # Unique
+General Inst. # Avg
+Steps
+RICO 72K - - -
+RICO semantics 72K - - -
+VINS 4K - - -
+MUD 18K - - -
+PixelHelp 800 - 187 4.2
+UGIF 3.3K - 480 6.3
+MoTIF 21K - 480 4.5
+AitW * 510K - 1539 6.5
+AitZ * $\blacksquare$ 18K 18K 2504 7.5
+AndroidControl * 99K - 15,283 4.8
+AMEX 104K 296K 3046 12.8
   • * indicates elements are mis-annotated and the bounding boxes are not well-aligned.
   • $\blacksquare$ indicates containing only action results of the element interacted at each step.
-However, the latest GUI agents, such as CogAgent [10] and AppAgent [28], exhibit unsatisfactory performance when handling real-world tasks.
-The primary issues with these agents are their lack of awareness of page layouts and their limited understanding of the functionalities of various user interface elements.
-These shortcomings are largely due to the absence of a comprehensive, detailed, large-scale dataset.
-Recent research has introduced several related datasets for the Android operating system, featuring different types of annotations.
-For instance, AitW [19] is a large-scale dataset annotated with instructional operations and screenshot icon detection.
-However, the data quality, especially in element-level annotation which is constructed by a pretrained IconNet, is subpar due to the unregulated annotation style.
-Besides, only 2.8% of instructions are for general third-party apps, not representative of general usage.
-Additionally, AitW suffers from heavy data redundancy as 94.9% instructions, including both multi-step and single-step, are in the category of “Web-shopping”.
-To address these issues, AitZ [32] selects a smaller subset of AitW and re-annotates it with screenshot descriptions and Chain-of-Action-Thought (CoAT) annotations.
-Though the CoAT paradigm is of some help, the dataset scale shrinks substantially to total 18K screen-action annotation pairs.
-Other datasets [2, 14, 23] also have various limitations, including the number and practicability of instructions, diversity, accuracy and reliability of annotations, coverage of apps and Android versions.
-Furthermore, these datasets primarily offer supplemental tree-based representations, such as the View Hierarchy.
-Notably, the View Hierarchy is not universally available across all apps, nor is it readily accessible to average mobile users.
-This limitation restricts the practical usability and generalizability of the data, as it does not reflect the typical interaction scenarios encountered by most users.
-When human users receive an instruction, they perform several steps beyond merely interpreting the instruction.
-A human user would comprehend the overall page layout and analyze the current page to identify interactive elements, determining which elements can be tapped and where to scroll.
-The functionality of those elements are then assessed, particularly the clickable ones.
-For example, a bell icon likely indicates navigation to a notification page, while a magnifier icon might signify a search page or activate a search bar.
-Equipped with this information and contextual knowledge, the user can effectively break down the instruction into specific actions for multiple screens.
-To align an GUI agent with humans, we design a comprehensive multi-level annotation dataset for the Android OS, named the Android Multi-annotation EXpo (AMEX).
-AMEX consists of three levels of annotations (see Figure 1): (i) GUI interactive element grounding, (ii) GUI screen and element functionality descriptions, and (iii) instructions with GUI-action chains.
-In total, AMEX consists of annotations on over 104K high-resolution screenshots, 711K element-wise functionalities, and around 3,000 unique instructions with stepwise GUI-action chains, with an average of 13 steps per instruction (see Table 1 and Table 2).
-The bounding boxes of GUI interactive elements are filtered and verified by human annotators.
-Screen descriptions and element functionalities, generated by GPT with built-in app descriptions, are also manually checked.
-Given random apps and instructions, annotators, all trained with specific and precise guidelines for each action, conduct the operations in a natural manner to achieve the objective.
-The data is collected from 110 third-party apps with Android 13 on Pixel 7 Pro and Samsung S10 emulators, ensuring up-to-date and prevalent screen resolutions and aspect ratios.
-The key features of AMEX can be summarized as providing knowledge of the modern mobile GUI environment from multiple levels, and complex human logics and operations on third-party apps.
-Our contributions can be summarized as follows: (a) We collect and release AMEX for training and evaluating generalist mobile GUI agents, which has multi-level annotations, providing reliable understandings and complex instructions of the smartphone UI environment; (b) We release the SOTA agent SPHINX Agent, which can serve as the baseline model for future researches on GUI agents.
-2 Related Work
-GUI-Control Datasets.
-Table 1 compares several popular GUI-control datasets on Android.
-While some works [5, 15, 20] focus on the web platform, on Android OS, many works [19, 29, 33] have focused on identifying various types of GUI elements, often assigning numerous classes to different elements.
-Other studies [2, 14, 23] primarily emphasize action-observation pairs during instructional operations, but their annotations are limited and often require supplemental View Hierarchy (VH) data for each screenshot.
-Furthermore, these instructions are typically too simplistic for real-world tasks.
-AitW [19] provides both screen GUI element annotations and instructions, but it includes only a small portion of instructions for third-party apps, with most operations conducted on Chrome and other system-built apps.
+Table 1: Comparison of mobile GUI-related datasets.
+Scale: the number of unique instructions on general third-party apps, average steps per instruction, number of screenshots and element functionality descriptions.
+Diversity: screenshot description, element labels, element functionality and the action details for stepwise operation.
+Table 2: AMEX statistics
+# Screenshots # Apps # Interactive Elements # Functionalities # Instructions # Avg. Steps
+104,876 192 1,659,647 296,075 3046 12.8
+Table 1 compares several popular GUI-related datasets on Android.
+While some works Deng et al.
+(2024); Liu et al.
+(2018); Shi et al.
+(2017); Wu et al.
+(2023) focus on the web or desktop platforms, on Android OS, many works Deka et al.
+(2017); Li et al.
+(2020); Feng et al.
+(2024); Bunian et al.
+(2021) have focused on identifying various types of GUI elements, assigning classes to different elements.
+This annotation style can be defined as a traditional icon classification and detection problem.
+In many cases, however, the functionality of an icon depends on the context of the GUI and many elements on mobile devices are not an icon and their functionalities are also critical.
+Besides, most of the datasets Deka et al.
+(2017); Bunian et al.
+(2021); Li et al.
+(2020) are outdated, where the screenshot interfaces are completely different from modern apps.
+The most recent MUD Feng et al.
+(2024) only provides 18K screenshots, which are insufficient for understanding the layouts and elements for LLMs.
+Other studies Burns et al.
+(2021); Rawles et al.
+(2024b); Zhang et al.
+(2024); Gubbi Venkatesh et al.
+(2024) primarily emphasize action-observation pairs during instructional operations, but their annotations are limited and often require supplemental View Hierarchy (VH) data for each screenshot.
+AitW Rawles et al.
+(2024b) provides both screen GUI element annotations and instructions, but it includes only a small portion of instructions for third-party apps, with most operations conducted on Chrome and other system-built apps.
 Additionally, each instruction is repeated multiple times, resulting in significant data redundancy.
-It also relies on the pre-trained IconNet for automatic annotations, which unfortunately leads to numerous mis-annotated types and misaligned bounding boxes.
-In response, AitZ [32] filters AitW thoroughly, selecting 2.5K unique instructions and episodes, and introduces the Chain-of-Action-Thought framework to annotate action results and page descriptions better.
+In response, AitZ Zhang et al.
+(2024) filters AitW thoroughly, selecting 2.5K unique instructions and episodes, and introduces the Chain-of-Action-Thought framework to annotate action results and page descriptions better.
 Despite this refinement, AitZ contains only 18K screen-action pairs.
-GUI-Control Agents.
-Recent advancements have leveraged the extensive world knowledge and robust embodied capabilities of Large Language Models (LLMs) [1, 8, 9, 11, 12, 22] for complex task planning and reasoning within GUIs.
-A notable approach involves employing business-level generalist models like GPT-4v directly as GUI-control agents.
-Works [28, 34] have utilized these models, employing extensive prompt engineering to guide the LLM in executing complex tasks.
-However, the effectiveness of these methods is inherently limited by the capabilities of the available generalist MLLMs [4, 13, 16, 24] and requires meticulous prompt design to achieve optimal results.
-Alternatively, another research line focuses on fine-tuning smaller LLMs on GUI-specific datasets to imbue them with domain-specific knowledge, thereby enhancing their operational efficiency.
-For example, CogAgent [10] enhances performance in GUI-related tasks by integrating a high-resolution cross-module that fuses image features from various levels.
-Similarly, MobileAgent [6] improves sample handling and structuring of input data to make it more consistent and LLM-friendly.
-Auto-UI [30] utilizes a multimodal encoder-decoder language model based on T5 and BLIP and CoCo-Agent [17] takes element OCR layouts as additional input to enhance the performance.
-Expanding beyond traditional mobile phone GUI domains, recent studies [26, 31] are exploring the potential of LLM-Agents in operating system-level task execution.
-3 Android Multi-annotation EXpo (AMEX)
+AndroidControl Li et al.
+(2024) is another large-scale instruction-based dataset, however, the element-wise annotations contain heavy redundancy and misaligned cases due to the lack of human verification.
+Figure 2: Overview of the data collection pipeline.
+The raw data is from two subsets collected by human annotators and an autonomous tool.
+Annotators record the GUI-action chains simultaneously while collecting the screenshots and corresponding XMLs.
+Then raw data is sent to annotators to filter GUI element bounding boxes, and then the boxes with corresponding raw screenshots are sent as input to GPT and Gemini to extract the GUI screen and element descriptions, which are then manually checked by humans.
+2.2 GUI-related Systems and Agents
+Wang et al.
+(2023) first starts to apply Large Language Model (LLM) on GUI, but it remains on tasks only interacted on single page, which are more like question-answer tasks rather than end-to-end instructional tasks.
+Recent advancements have leveraged the extensive world knowledge and robust embodied capabilities of LLMs Gu and Dao (2024); Gur et al.
+(2023); Touvron et al.
+(2023) for task planning and reasoning.
+Works Zhang et al.
+(2023); Zheng et al.
+(2024) have utilized business-level models (e.g., GPT-4V), employing extensive prompt engineering to guide the LLM in executing complex tasks.
+The effectiveness of these methods requires a meticulous prompt design to achieve optimal results.
+Alternatively, another research line focuses on fine-tuning smaller LLMs on GUI-specific datasets to imbue them with domain-specific knowledge.
+For example, CogAgent Hong et al.
+(2024) enhances performance in GUI-related tasks by integrating a high-resolution cross-module that fuses image features from various levels.
+CoCo-Agent Ma et al.
+(2024) and AndroidControl Li et al.
+(2024), unlike other agents taking only screenshots as input, use element layouts from accessibility trees or view hierarchy as additional input to enhance the performance.
+However, still many apps don’t support accessibility information or only provide very little of that.
 (a) Element annotation on AitW (left) and AMEX (right).
@@
 Yellow box: vertically scrollable element.
-Figure 3: Overview of the data collection pipeline.
-The raw data is from two subsets collected by human annotators and an autonomous tool respectively, while annotators record the GUI-action chains simultaneously.
-Then raw data is sent to annotators to filter GUI element bounding boxes, then they with raw screenshots are sent as input to GPT to extract the GUI screen and element descriptions, which are then manually checked by humans.
+3 Android Multi-annotation EXpo (AMEX)
 When receiving an instruction, a human user first analyzes the overall screen layout to form a basic understanding of the current Android environment.
@@
 Based on this human cognitive process, we design three levels of annotations in our proposed AMEX training set: (i) GUI interactive element grounding (see Section 3.2), (ii) GUI screen and element descriptions (see Section 3.3), and (iii) instructions with GUI-action chains (see Section 3.4).
-The raw data is collected using Android emulators, specifically Android Virtual Device and Genymotion [^1] https://www.genymotion.com/.
-We developed tools to autonomously perform emulator operations as well as record human actions.
-These tools leverage Appium [^2] https://appium.io/docs/en/latest/, an open-source, cross-platform test automation tool for Android, iOS, and web apps.
-Appium collects device screenshots during operations, each of which also corresponds to an Extensible Markup Language (XML) file recording the basic screen layout with element attributes, such as bounding boxes and in-app descriptions.
+The overall statistics of AMEX are listed in Table 2.
 3.1 Data Collection Pipeline
@@
 The raw data is collected through two methods: human instruction-following GUI manipulations and autonomous GUI controls.
-Human GUI manipulations involve recording stepwise operations for each instruction, and simultaneously storing screenshots and each screen’s XML data before each stepwise operation.
+Human GUI manipulations involve recording stepwise operations for each instruction, and simultaneously storing screenshots and each screen’s Extensible Markup Language (XML) data.
 In parallel, an autonomous script controls emulators to collect additional screenshots and their XMLs.
@@
 Human annotators then review each screenshot to filter out all the misaligned boxes, which serve as the interactive element grounding annotations.
-With the in-app descriptions, GPT generates the functionalities of the selected elements and provides descriptions for the whole screenshot.
-Human annotators then further check the descriptions of functionalities.
-More detailed methods are discussed in the following sections and Appendix A.1.3.
+With the in-app descriptions, GPT and Gemini generates the functionalities of the selected elements and provides descriptions for the whole screenshot.
+Annotators then further check the quality of the descriptions of functionalities.
+More details are discussed in the following sections and Appendix A.1.
 3.2 Level I: GUI Interactive Element Grounding
-Existing datasets [19, 32] typically classify elements on the screen, such as icons, texts, and images, based on their types.
+Existing datasets Deka et al.
+(2017); Sunkara et al.
+(2022); Feng et al.
+(2024); Bunian et al.
+(2021) typically classify elements on the screen, such as icons, texts and images, based on their types.
 Instead of adhering to the traditional classification paradigm, we define interactive elements more broadly as any elements that users can interact with, regardless of their specific types (see Figure 3(a)).
 Specifically, interactive elements in our dataset are only categorized into two subsets: (i) clickable elements and (ii) scrollable elements.
+The motivation of our annotation is mainly from two facts: (i) no existing dataset provides information on the “scroll” action, which is crucial especially for agents to know the horizontally scrollable area.
+(ii) “click” action is the most used action and many “compound” elements cannot be simply categorized.
+As illustrated in Figure 3(b) red boxes, four “compound” elements in the middle contain both image and text and the definition of “clickable” is more suitable with the functionality in Level 2 for agents to understand the element.
 Clickable Elements are the most common components in a screen.
 They typically include clickable icons, images, texts, and compounds that combine several categories.
-Figure 2(b) illustrates various clickable elements in red boxes.
 We also include certain “typeable” elements, such as search bars, because most typeable elements require a prior click action to enable typing.
 Scrollable Elements typically occupy larger areas on the screen.
-In most cases, a scrollable element area supports a pair of actions, such as “scroll down” and “scroll up” or “scroll left” and “scroll right.” Generally, the vertical scrollable area forms the main frame in the middle of the screen, while the horizontal scrollable area serves as a “carousel” for listing different categories or activities.
-Figure 2(b) illustrates these two types of scrollable elements.
-As mentioned in Section 3.1, the collection of screenshots and their XML data utilizes two methods.
-Here, we provide a more detailed pipeline for executing an autonomous script to perform operations on an Android emulator.
-The script is designed to traverse an app in an unconstrained manner and collect data.
-It performs actions (see Appendix A.1.1) at regular time intervals to allow each page to fully load.
-After waiting, the script captures the current screenshot along with the corresponding XML file.
-To ensure efficiency and prevent redundancy, we limit the maximum number of operation steps per script execution.
-Additionally, a break mechanism is implemented to exit potential dead loops, ensuring the script progresses effectively.
-Upon the raw data (yellow part in Figure 3), we extract bounding boxes for interactive elements from the XML associated with each screenshot.
-While scrollable elements are typically parsed with high quality, clickable elements are more problematic.
-This issue is from the fact that app developers often design elements with overlapping or layered coverage, leading to the generation of bounding boxes even for elements that are visually hidden from the user’s view.
-To address this, annotators manually inspect each screenshot, identifying and selecting only the interactive elements that are actually visible within the interface (see Appendix A.1.2).
+In most cases, a scrollable element area contains many clickable elements and supports a pair of actions, such as “scroll down” and “scroll up” or “scroll left” and “scroll right.” Figure 3(b) illustrates these two types of scrollable elements.
 3.3 Level II: GUI Screen and Element Functionality Descriptions
-Previous works on GUI elements often rely on predefined class names to convey the underlying meaning of each element.
+Previous works Deka et al.
+(2017); Feng et al.
+(2024); Bunian et al.
+(2021) on GUI elements often rely on predefined class names, such as “image” and “text”, to convey the underlying meaning of each element.
+Rico Semantics Sunkara et al.
+(2022) annotates the icons with slightly more detailed semantics information but the annotations are still classifications on icons.
 However, this classification-based method has significant limitations.
-For instance, a small triangle symbol typically indicates “play” in video or music interfaces, but it might also appear as part of a content creator’s profile image.
-In such cases, the model erroneously interprets the symbol and repeatedly attempts to tap the icon to play the video.
-This issue arises because the class-based annotation approach focuses on class labels rather than truly understanding the functionality of each element in the surrounding context.
+For instance, a “plus” symbol typically indicates “plus” in calculator interfaces, but it means “create a new task” in a ToDo interface.
+This class-based annotation approach focuses on class labels rather than truly understanding the functionality of each element in the surrounding context, which may lead to a misunderstanding of the page layout.
 To ensure the dataset is truly instructive rather than merely icon detection, we focus on describing screen status and element functionalities.
-Consider the above mentioned example: instead of providing a bounding box for the small triangle and labeling it as “ICON_PLAY,” we present the actual functionality of the element within its context (e.g., “Opens the creator’s personal page”).
-This strategy offers clear, instructive, and detailed description, enhancing the dataset’s applicability.
-We adopt GPT-4o for the element-wise annotation to describe screens and interpret element functionalities.
-Given the compact and dense nature of GUI elements, we apply Set of Markers (SoM) techniques [27] to bolster GPT-4o’s capability for visual localization.
-The coordinates of these elements are obtained and cleaned in the preceding phases (Section 3.2).
-Additionally, recognizing that some elements possess abstract functionalities that are challenging to discern, we supplement the prompts for GPT-4o with in-app descriptions (if available) extracted from XMLs to enhance GPT-4o’s comprehension of the screen.
-After the data generation, human annotators manually check 20% of the descriptions of the element functionality.
-Further elaboration on this methodology is provided in Appendix A.1.3.
+Consider the above mentioned example: instead of providing a bounding box for the icon and labeling it as “ICON_PLUS”, we present the actual functionality of the element within its context (e.g., “Create a new task”).
+This strategy offers a clear, instructive, and detailed description of the interactive element, enhancing the dataset’s applicability.
+Besides, the traditional categories of “IMAGE” and “TEXT” provides very little information of the actual element functionality.
+Our annotation will provide a much more detailed and informative functionality for those elements (e.g., “View the detail page of the ’Jazz Classics’ playlist” for the red left bottom “compound” element in Figure 3(b)).
+Further elaboration on the collection and processing is provided in Appendix A.1.4.
 3.4 Level III: Instructions with GUI-Action Chains
-Our focus for GUI agents is predominantly on third-party applications rather than system-built apps.
-We select an average of three apps per category from the Google Play Store.
-Due to the significant increase in app functionalities over the past years, we estimate that a sample size of approximately 50 apps can adequately represent most common use scenarios.
-The instruction generation process comprises three phases.
-Initially, human annotators create 5-10 complex instructions for each target app, considering its specific purposes and capabilities.
-These initial instructions, combined with relevant app metadata collected online, serve as input for ChatGPT.
-It then generates a larger set of 80-100 instructions that exhibit similar structure and intent to the human-provided examples.
-For example, if a human annotator provides the instruction “Open Google Maps.
-Find the shortest route from the current location to Empire State Building by car drive,” ChatGPT might produce “Open Google Maps.
-Find the fastest route from the current location to Rockefeller Center by public transit.” However, due to ChatGPT’s limitations on understanding real-world constraints, human filtering is adopted to remove any unreasonable or impractical instructions.
-Notably, instructions within AMEX are more complex than those in the AitW dataset, with an average step count of 12.8, nearly double that of AitW.
+Table 1 illustrates that most instruction-based datasets provide an average number of operation steps below 7, which implies most of the instructions are simpler than tasks in real-world.
+For example, users often apply filters or sort by different attributes when searching, which causes more complex instructions and longer action chains.
+In order to address this situation, we collect a set of complex instructions with GUI-action chains on popular apps from Google play store.
 We define our action space for stepwise GUI operations similarly to AitW: {TAP, SCROLL, TYPE, PRESS_BACK, PRESS_HOME, PRESS_ENTER, TASK_COMPLETE, TASK_IMPOSSIBLE}.
@@
 Additionally, for instructions that involve information query (e.g., “What is the lowest price of the men’s belt?”), we associate the TASK_COMPLETE action with a region of interest, which is defined as the bounding box of the area on the current screenshot, where the answer is expected to appear (see examples in Appendix A.3).
-This comprehensive action space allows us to fully simulate a wide range of typical use cases.
-The collection tool is designed to simultaneously record the screen and action pair together with the XML file. The details of GUI-action chain collection can be found in Appendix A.1.4.
-3.5 AMEX Statistics and Test Set
-Table 2: AMEX statistics
-# Screenshots # Interactive
-Elements # Functionalities # Instructions # Avg. Steps # Apps
-103,684 1,658,439 711,823 2,946 12.8 110
-Figure 4: (a) The number of screenshots by the number of GUI clickable elements. (b) The number of instructions by the number of GUI operations steps. (c) The word cloud of the instructions.
-Data statistics.
-Table 2 lists the statistics of the AMEX dataset.
-AMEX has approximately 104K screenshots, 1.6M GUI interactive elements for grounding, 712K GUI element functionality descriptions, and 3K instructions.
-Figure 4 illustrates the distribution of GUI interactive elements per screen and GUI operation steps along with the word cloud of the instructions.
-Test set.
-Instructions and their annotated screen-action chains from 9 out of 55 apps are chosen as the test set, including those on Booking, Gmail, NBC News (News), SHEIN, Citymapper (CM), Microsoft To Do (ToDo), Signal, Yelp and YouTube Music (Music), covering different scenarios.
-The number of test instructions is 362, around 10% of the total instructions.
-The evaluation of AMEX test set is similar to the existing dataset AITW [19].
-We perform the action matching to get the partial action match score.
-The partial action matching score is the number of correct actions divided by the episode length [14].
-To ensure the correctness of the test set, we carefully review the trajectory annotations in the unseen test set, where we drop the former step of PRESS_BACK, which indicates the wrong operation of the annotator.
-However, we keep the PRESS_BACK step to evaluate the agent’s self-correction ability.
-4 SPHINX-GUI Agent
-Integrating data from AMEX with the existing AitW dataset, we train the SPHINX-GUI Agent tailored for GUI-related tasks as the baseline for AMEX dataset.
-This Agent is initialized from SPHINX [7] and is specifically trained to generate actions to manipulate the GUI effectively, which are different from the more flexible output formats typical of standard MLLM tasks such as QA.
-To incorporate GUI-specific intelligence into SPHINX, we have implemented several key modifications:
-Instruction Dataset Construction.
-We transform the AMEX data and the publicly available AitW dataset into a VQA format conducive to instruction understanding, thereby adapting the MLLM as a GUI agent.
-To embed GUI-specific knowledge within the MLLM, we developed four distinct VQA tasks, detailed in the instruction templates provided in Table 3.
-Table 3: Task template examples.
-Tasks Examples of Task Templates Source
-Screen Description
-User: Provide a one-sentence caption for the provided GUI image.
-Agent: This page is a film section of the IMDb app, displaying a list of movies or shows with sorting and filtering options.
-AMEX
-Element Grounding
-User: Identify all clickable elements and provide their 2D bounding boxes.
-Agent: List of Elements with BBoxes.
-AMEX
-Functionality Description
-User: What is the function of the icon at the coordinates: BBox.
-Agent: Click to expand or collapse the sort options.
-AMEX
-GUI-Action Prediction
-User: Please execute the task described by the following instruction: Delete the first email in the inbox.
-Agent: CLICK[[0.59, 0.92]].
-AMEX & AitW
-Structured Representation.
-To effectively guide the agent in task completion, we incorporate the goal into the input prompt.
-To prevent redundant actions, we also include a history of actions formatted as $X_{\text{history}}=\left[a_{t-k},\ldots,a_{t}\right]$, where each $a$ denotes an action type, accompanied by a touch point for actions such as SCROLL or TAP.
-The parameter $k$ denotes the length of the history.
-Action trajectories from both AitW and AMEX datasets are converted into this uniform representation.
-Given the extensive nature of GUI screenshots and the small relative size of some target elements within the screen, we adopt the “any resolution” approach from SPHINX [7].
-This method initially partitions the input image into sub-images, then encodes separately.
-The LLM processes all the resulting visual tokens to ensure comprehensive understanding and interaction capabilities.
-We follow all other default training receipts in [7] to obtain the final agent.
+This comprehensive action space allows users to fully simulate a wide range of typical use cases.
+Detailed collection process and instruction generation are in Appendix A.1.5.
 4 Experiments
-5.1 Evaluation Setup and Compared LVLMs
-We evaluate different GUI agents including SPHINX Agent developed in Section 4 on AitW and our AMEX dataset respectively.
-For AitW, we simultaneously train SPHINX Agent on all the subsets and then assess all test sets.
-We downsample GoogleApps subset to $10\%$ to avoid data imbalance.
-The training and test split setting follows the one in [19].
-For AMEX, we evaluated our SPHINX agents on the test set as described in Section 3.5.
-Our models are benchmarked against several agents including Auto-UI [30], SeeClick [3], CogAgent [10] and CoCo-Agent [17].
-5.2 Experiment Results
-The experiment outcomes for AitW and AMEX are presented in Tables 4 and 5 respectively.
-As indicated in Table 4, SPHINX Agent (SphAgent) trained on AitW exhibit competitive performance on the AitW dataset relative to the baselines.
-Notably, the introduction of the AMEX dataset enhanced the overall performance by approximately 2.5%, with significant gains observed in the “General” (5%), “Single” (5%) and “WebShopping” (2%) tasks, which indicates the strong complement from AMEX dataset.
-The GUI element functionality descriptions are served as the “Single” step instructions which strongly promote the performance in the “Single” tasks, while 3K instructions from general third-party apps also boost the performance in the “General” tasks.
-“WebShopping” is the category with the most complicated tasks in AitW, which also benefits from the complex instructions provided by AMEX.
-Table 4: Experiment results on AitW. + indicates CoCo-Agent takes screen element OCR layouts as additional inputs.
-Agent # Params Training Data General Install G-Apps Single WebShopping Overall
-Auto-UI 5B AitW 68.24 76.89 71.37 84.58 70.26 74.27
-[…]
-[…]
-CoCo-Agent+ 7B AitW 70.96 81.46 76.45 91.41 75.00 79.05
-SphAgent 7B AitW 68.2 80.5 73.3 85.4 74 76.28
-SphAgent 7B AitW + AMEX 73.1 80.5 73.4 90.8 75.8 78.72
-Table 5: Experiment results on AMEX
-Agent Training Data Gmail Booking Music SHEIN News CM ToDo Signal Yelp Overall
-[…]
-SphAgent AitW 32.1 45.9 46.1 35.1 48.3 61.1 55.9 43.3 42.9 45.63
-SphAgent AMEX 61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.71
-SphAgent AitW + AMEX 62.4 68.1 76.3 71.9 68.6 67.3 77.6 66.0 64.1 69.14
+AMEX serves as a supplementary dataset for other large-scale instruction-based datasets.
+To test the effectiveness of AMEX, we directly finetune SPHINX models Liu et al.
+(2024) without any tricks and conduct evaluation on three other benchmark datasets.
+We choose ScreenSpot Cheng et al.
+(2024) as the benchmark of our agent’s element grounding ability and AitW and AndroidControl as the benchmarks of our agent’s GUI-control ability.
+4.1 SPHINX Agent
+Our SPHINX agent, SphAgent for short, is initialized from SPHINX Liu et al.
+(2024) and it is finetuned for the GUI-control tasks and element grounding tasks.
+For both tasks, we train the model in a pure vision-based principle, which means the model understands the environment by only screenshots, without any other supplementary information such as accessibility trees or view hierarchy, due to the fact that many apps don’t support or only provide very little accessibility tree or view hierarchy element information.
+We consider vision-based agents are more applicable in more various use cases.
+Table 3: ScreenSpot mobile subset experiment results.
+Model Size Icon / Widget
+Fuyu 8B 1.3%
+CogAgent 18B 24.0%
+SeeClick 9.6B 52.0%
+Qwen2-VL 7B 60.7%
+GPT-4V w. OmniParser - 57%
+SphAgent 7B 72.6%
+4.2 ScreenSpot
+Table 4: Experiment results on AndroidControl. “Click & Long Press” is specially computed to show the improvements on these actions. The best results are in bold for different subsets and levels.
+Agent # data Training Data Task
+Level IDD Category
+Unseen App
+Unseen
+Task
+Unseen Overall Click &
+Long Press
+SphAgent 178K AndroidCtrl High Lv. 58.3 40.6 39.4 52.2 49.8 34.1
+Low Lv. 79.8 71.0 71.1 83.5 75.8 53.0
+SphAgent 178K 70% AndrCtrl High Lv. 60.4 41.7 43.4 59.3 52.8 37.3
++ 10% AMEX Low Lv. 81.9 73.2 72.6 87.5 77.7 56.7
+SphAgent 712K AndroidCtrl High Lv. 70.5 51.6 55.6 70.2 61.7 50.8
++ AMEX Low Lv. 88.9 81.8 76.9 92.6 84.2 67.0
+ScreenSpot Cheng et al.
+(2024) is a benchmark dataset which contains over 600 UI screenshots from mobile devices, desktops and websites.
+It is designed to evaluate the model’s element grounding capability, where each human-annotated functionality corresponds to an interactive element.
+We train SPHINX agent on AMEX level 1 and level 2 data to evaluate the element grounding capability.
+Table 3 lists the evaluation results on the ScreenSpot mobile subset.
+The models Bavishi et al.
+(2023); Hong et al.
+(2024); Cheng et al.
+(2024); Wang et al.
+(2024) in first four rows are LVLMs specially trained on GUI data.
+OmniParser Wan et al.
+(2024) is a combination of an icon detection model, an icon description model and an OCR module, which serves as a screenshot parser for other LVLMs such as GPT-4V.
+During evaluation, our SphAgent surpasses four LVLMs by a large margin on the “Icon / Widget” subset without any training tricks, proving that the functionality understanding can largely boost the performance of GUI grounding for agents.
+The above results and comparisons strongly prove the effectiveness and applicability of AMEX in GUI element grounding tasks.
+4.3 AndroidControl
+AndroidControl Li et al.
+(2024) is a large-scale benchmark dataset to evaluate the agent’s GUI-control ability to complete tasks on general apps (see more detailed description of the dataset and splits in Appendix A.2.1).
+We evaluate three SphAgents trained on (i) only AndroidControl, (ii) random 70% AndroidControl data mixed with random 10% AMEX level 1 and 2, (iii) a mix of AndroidControl and AMEX level 1 and 2.
+The (ii) experiment ensures the same number of data point as in (i) experiment.
+The action space of AndroidControl is different from AMEX level 3 and AitW, so we only use level 1 and level 2 data during the experiments (ii) and (iii).
+Table 4 lists the evaluation results on AndroidControl test set.
+The evaluation results of experiment (i) and (iii) strongly support the effectiveness of AMEX level 1 and level 2.
+Adding full data of AMEX level 1 and level 2 leads to an average 10% overall performance gain and strongly improves the “click” and “long press” actions by more than 14%.
+Also, the comparative results from experiment (i) and (ii) shows that even using the same number of data points, replacing down-stream instructional task data with environment understanding data would also lift the performance of agents on both low-level and high-level at an average of 2.5%.
+This performance gain indicates the effectiveness of our multi-level annotations.
+4.4 AitW
+AitW is another large-scale benchmark dataset to evaluate the agent’s GUI-control ability to complete task goals.
+We evaluate three SphAgents trained on (i) only AitW, (ii) a mix of AitW and AMEX.
+Since the GUI-action chains in AMEX share the same action space with AitW, we utilize all three levels of AMEX in the mixed data.
+Table 5 lists the evaluation results on AitW test set, which has five subsets for different types of tasks.
+SphAgent trained on the mixed data (AitW + AMEX) achieves the highest overall accuracy among all agents, and when compared to the SphAgent trained on only AitW, the accuracies in “General” and “Single” categories have a notable gain of around 5% and the overall accuracy lifts by 2.5%.
+That the gains in other categories are not remarkable is possibly due to the mis-aligned element-wise annotation and unregulated instruction operation annotation.
+In Appendix A.2.2 we further explore the dataset evaluation and explain the reason why the evaluation results might be misleading.
+Table 5: Experiment results on AitW.
+Agent Training Data General Install G-Apps Single WebShopping Overall
+SphAgent AitW 68.2 80.5 73.3 85.4 74 76.28
+SphAgent AitW + AMEX 73.1 80.6 73.4 90.8 75.8 78.72
+SphAgent AMEX 51.5 55.6 57.1 61.9 55.1 56.2
+Table 6: Experiment results on AMEX. The test apps are excluded during the training of SphAgent on AMEX.
+Agent Training Data Gmail Booking YT-Music SHEIN NBC CM ToDo Signal Yelp Overall
+SphAgent AitW 32.1 45.9 46.1 35.1 48.3 61.1 55.9 43.3 42.9 45.6
+SphAgent AMEX 63.7 67.4 76.1 70.0 68.5 62.7 78.6 68.2 68.9 69.3
 4.5 Cross-Domain Experiments
-We also examine the generalization capabilities of our agent across different domains.
-Specifically, the agent trained exclusively on the AitW dataset is tested on the AMEX dataset.
-The results, as detailed in the first row of Table 5, reveal significant findings.
-Agent trained only on AitW, shows satisfactory performance on the AitW dataset itself.
-However, when evaluated on the AMEX dataset, there is a notable decrease in performance.
-This performance degradation can likely be attributed to several factors: 1) Higher complexity in AMEX.
-The AMEX dataset generally involves longer operational sequences compared to those in AitW, presenting a higher complexity level that the agent might not have been exposed to during training.
-2) Domain gap.
-There are substantial differences in the visual textures and language instructions between the two datasets.
-These variations could hinder the agent’s ability to effectively transfer its learned knowledge to a new, unfamiliar domain.
-Additionally, we observed a significant performance degradation for the SeeClick agent, originally trained on AitW.
-Beyond the previously mentioned reasons, this decline may also stem from the fact that the original SeeClick model was designed to process only $224\times 224$ image inputs.
-In contrast, the AMEX dataset comprises high-resolution images and requires interactions with tiny icons.
-The high-definition images in AMEX introduce new challenges for GUI agents.
-5.4 AitW Human Evaluation Results
-Table 6: Human evaluation results on 5%-subsampled AitW test sub-set. “MMR” stands for “Mis-Match Random” subset.
-Subset Action Sources General Install G-Apps Single WebShopping Overall
-MMR SphAgent 90.91 83.65 87.50 87.68 82.39 86.43
-AitW Anno 88.18 75.00 82.69 84.78 75.57 81.25
-Random SphAgent 92.31 93.75 94.33 94.83 94.55 93.95
-AitW Anno 94.23 90.34 93.62 93.97 95.00 93.43
-As specified in AitZ [32], several types of error cases are identified in the AitW test set (see Figure5).
-To assess the unreliability of the original AitW test set and its evaluation methods, we propose AitW-HE (Human Evaluation), a refined subset of the AitW test set.
-Human annotators evaluate two subsets derived from the original test set.
-One subset is randomly chosen from episodes where SphAgent receives low scores (Mis-Match Random), i.e., about 66 score compared to the overall 78.72 shown in Table 4, indicating a high mismatch between inference results and AitW annotations.
-The other subset is randomly selected from the remaining episodes (Random).
-Annotators first filter out repeated and redundant screenshots (see Figure 5(a)).
-For each remaining screenshot, they then record whether the SphAgent-inferred action and the original AitW annotation are correct.
-Figure 6 illustrates cases where annotators mark both the inferred action and the original annotation as correct, even though they interact with different elements.
-(a) Repeating useless screenshots and actions.
-(b) Wrong task complete.
-Figure 5: Error cases in AitW test set. Red dots indicate the actions from the AitW annotations.
-Table 6 presents the accuracy scores from human evaluation.
-The table indicates that the low SphAgent scores in the MMR subset are primarily due to unsatisfactory annotations (AitW Anno), which are used as ground truth during the evaluation.
-Both SphAgent and AitW Anno in the “MMR” subset have lower scores than those in the “Random” subset, highlighting that the tasks in the MMR subset are relatively more challenging.
-Furthermore, SphAgent achieves better human evaluation results than the original annotations, demonstrating its effectiveness and close alignment with human judgment.
-Comparing the results from Table 4 and Table 6, the notable differences in overall scores (i.e., 78.72%, 86.43%, and 93.95%) underscore the unreliability and misleading nature of the AitW test set and its evaluation methods.
-Figure 6: Purple dots indicate the SphAgent inferred actions and red dots indicate the AitW annotations. Human annotators mark them both correct even though they are clicking different elements.
-5.5 Ablation Study on Multi-level Annotations
-To validate the effectiveness of the multi-level annotations proposed in the AMEX dataset, we conducted an ablation study.
-As shown in Table 7, the agent achieved a performance gain of approximately 1.0 when aided by L1 annotations.
-When trained with L1 and L3 annotations, the agent exhibited a performance gain of about 1.6 compared to the baseline.
-The gain from L2 annotations is approximately 2.3.
-Furthermore, training the agent with the complete set of annotations resulted in a performance gain of approximately 3.5.
-The ablation study results demonstrate that each level of annotations in AMEX enhances the final agent’s performance.
-Table 7: Ablation study of each level of annotations.
-Tasks Gmail Booking Music SHEIN NBC CM ToDo Signal Yelp Avg
-L1 L2 L3
-45.9 64.5 74.4 71.8 70.3 67.4 79.3 64.9 66.3 67.2
-58.6 68.3 74.4 68.8 68.1 61.1 79.3 73.2 67.4 68.8
-60.3 66.1 77.2 70.2 72.0 64.0 78.1 72.3 65.7 69.5
-61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.7
+Since AMEX Level 3 shares the same action space only with AitW, we conducted two experiments: (i) training SphAgent on AMEX and testing it on AitW, and (ii) training SphAgent on AitW and testing it on AMEX.
+The results, as shown in the last row of Table 5, indicate poor performance when the agent is trained only on AMEX and tested on AitW.
+Similarly, Table 6 demonstrates poor performance when the agent is trained on AitW and tested on AMEX.
+This significant discrepancy in evaluation results is likely due to the domain gap between the two datasets.
+First, the styles of collecting GUI-action chains differ between the two datasets.
+For instance, annotators in AitW tend to directly search for apps, whereas annotators in AMEX prefer navigating through the app library to locate apps.
+This inconsistency in interaction strategies poses a challenge for the agent, which evaluates tasks based on static frames rather than dynamically adapting to different usage patterns.
+Second, although both datasets share the same action space, they differ in terms of visual appearance and versioning.
+The apps in AMEX are updated to their latest versions, while those in AitW are based on older versions.
+These versioning differences lead to changes in the visual appearance of GUI elements (e.g., color schemes, icons, or button designs), which can negatively impact the agent’s ability to recognize and interact with them.
 5 Discussions
 5.1 Limitations and Future Work
-Multi-lingual    Most existing datasets are limited to English, with UGIF [23] being a notable exception, as it includes instructions and screenshots in eight languages.
+Multi-lingual
+Most existing datasets are limited to English, with UGIF Gubbi Venkatesh et al.
+(2024) being a notable exception, as it includes instructions and screenshots in eight languages.
 The AMEX dataset contains a small number of screenshots in Chinese and Spanish, primarily due to strict registration and login requirements for Chinese apps and a lack of expertise in other languages.
 Future work should incorporate multi-lingual screenshots, functionalities, and instructions to create a more robust and comprehensive multi-lingual environment for GUI agents.
-Evaluation    The current evaluation method employed in both AMEX and AitW is straightforward: the model predicts the current action based on given instructions, previous operations, and screenshots.
-However, this evaluation approach has significant limitations in real-world scenarios.
-It fails to account for loading pages, waiting times for page loading, etc.
-An ideal evaluation would be an “Arena” (a mobile version of the environment in WebArena [35]) — an independent local virtual device equipped with Android and static background apps.
-This setup would ensure consistent results for the same queries, avoiding variations due to differing dates, prices, and other dynamic data.
+Online Evaluation
+Due to the limitation that SPHINX is not deployable, we encountered issues trying online evaluation such as AndroidWorld Rawles et al.
+(2024a).
+In future work, online dynamic evaluation is more accurate and simulate the real-world agent usage.
+Multi-Platform
+Android is an open-source OS and the community provides well-developed tools for interactions and emulations.
+iOS is another widely used mobile device OS but it is restricted and hard to run test on.
+In the future, a dataset extended to other OS or platforms would provide more general information.
 5.2 Ethical Considerations
@@
 A.1 Pipeline Details
-A.1.1 Autonomous script details
+The raw data is collected using Android emulators, specifically Android Virtual Device and Genymotion [^1] https://www.genymotion.com/.
+We developed tools to autonomously perform emulator operations and record human actions.
+These tools leverage Appium [^2] https://appium.io/docs/en/latest/, an open-source, cross-platform test automation tool for Android, iOS, and web apps, under Apache license.
+Appium collects device screenshots during operations, each of which also corresponds to an XML file recording the basic screen layout with element attributes, such as bounding boxes and in-app descriptions.
+A.1.1 Collection details
+As mentioned in Section 3.1, the collection of screenshots and their XML data utilizes two methods.
+Here, we provide a more detailed pipeline for executing an autonomous script to perform operations on an Android emulator.
+The script is designed to traverse an app in an unconstrained manner and collect data.
+It performs actions (see Appendix A.1.2) at regular time intervals to allow each page to fully load.
+Then the script captures the current screenshot along with the corresponding XML file.
+Upon the raw data (yellow part in Figure 2), we extract bounding boxes for interactive elements from the XML associated with each screenshot.
+Then annotators manually inspect each screenshot, identifying and selecting only the elements that are actually visible within the interface (see Appendix A.1.3 for detailed reasons).
+A.1.2 Autonomous script details
 The autonomous script controls the emulator using three actions: TAP, SCROLL, TYPE.
@@
 These phrases are primarily used in search scenarios.
-A.1.2 GUI clickable element filtering
+A.1.3 GUI clickable element filtering
 Figure 4: Demonstration of human annotator filtering. (a) before filtering. (b) after filtering.
@@
 Thus we need human annotators to filter out these blocked elements to get Figure 4(b).
-A.1.3 GUI screen and element functionality description collection details
-We apply the Set-of-Mark (SoM) [27] technique when using GPT-4o as the description generator.
+A.1.4 GUI screen and element functionality description collection details
+We adopt GPT-4o and Gemini 1.5 Pro to describe screens and interpret element functionalities separately.
+Given the compact and dense nature of GUI elements, we apply Set of Mark (SoM) techniques Yang et al.
+(2023) to boost GPT and Gemini’s capability for visual localization.
+The coordinates of these elements are obtained and cleaned in the preceding phases (Section 3.2).
+Additionally, recognizing that some elements possess abstract functionalities that are challenging to discern, we supplement the prompts for GPT and Gemini with in-app descriptions extracted from XMLs to enhance their comprehension of the screen.
+After the generation, we further utilize GPT-4o to compare the functionalities of the same element generated by two models and keep them if two functionalities have the same meaning.
+Finally, human annotators verify the quality and the functionality accuracy rate is above 97%.
+We apply the Set-of-Mark (SoM) technique when using GPT and Gemini as the description generator.
 The SoM technique is a visual prompting method designed to enhance the visual grounding capabilities of large multimodal models (LMMs), such as GPT-4V, by overlaying visual marks on image regions.
@@
 Figure 5 shows the screenshot with SoM technique.
+The most recurring mistake is the mis-interpretation of the content description.
+For instance, some “compound” elements, which may contain several text, rating stars, image, etc., have couple of content descriptions.
+Those mixed content descriptions may cause the misunderstanding of the real functionality of the compound element.
+In IMDB app, a row of movie entry sometimes would consist of a poster, a text name, ratings of the movie, and the price of renting.
+The desired functionality of the entry could be "view the detail page of the movie xxx", while GPT might generate "view the price for renting the movie xxx", which is not exactly what we want.
+However, we use two LLMs to double check their generations on the same element functionality and we only keep the elements with the same functionalities generated by two LLMs.
 Figure 5: Demonstrations of SoM technique on screenshots.
-The prompt is in the following format.
+The prompt is in the following Listing 1 format.
 Based on the screenshot of an Android mobile phone from the APP_NAME, please follow the instructions:
@@
 Listing 1: An example prompt for guiding GPT4o to generate the element functionalities for the given Screenshot.
-A.1.4 GUI-Action chain collection details
+A.1.5 GUI-Action chain collection details
+The instruction generation process comprises three phases.
+Initially, human annotators create 5-10 complex instructions for each target app, considering its specific functions and capabilities.
+These initial instructions, combined with relevant app metadata collected online, serve as input for GPT.
+It then generates a larger set of 80-100 instructions that exhibit similar structure and intent to the human-provided examples.
+However, due to GPT’s limitations on understanding real-world constraints, human filtering is adopted to modify or remove any unreasonable or impractical instructions.
+For example, the human-proposed instruction “Open Booking.com, search for stays in Washington DC from 2024 June 11 to June 15.” would generate a similar but impractical “Open Booking.com, search for stays in New York from 2024 May 3 to May 7”, which needs modification since the dates are not selectable when the task is generated.
+Since the meta information online doesn’t reflect all functions of an app and sometimes may include misleading information, the instruction generated by GPT may have two types of impracticabilities: (i) the task is unrelated to the app activity (ii) some of the steps in the instruction are impractical.
+For instance, an app “SEPHORA” only sells beauty related products, but GPT may generate tasks that searches for electronics.
+Another example would be about a ticket booking app “SeekGeek”, which doesn’t provide a sort option “sort by price from high to low”.
+While human created task would include “sort by price from low to high”, GPT would generate something with“sort by price from high to low”.
+This step is impractical so we would filter it out.
 Human annotators are each assigned a random selection of apps and their associated instructions, which they are asked to complete in a natural manner.
@@
 Additionally, after completing an information query task, annotators are asked to manually mark the region of interest on the screenshot using our annotation tool.
-A.1.5 Collection resources details
+A.1.6 Collection resources details
   • GUI interactive element grounding takes approximately 3000 human-hour to filter bounding boxes described in Appendix A.1.3.
-  • GUI screen and element descriptions use GPT-4o API, which consumes about 600 dollars.
-  • Instructions with GUI-action chains take approximately 200 human-hour.
+  • GUI screen and element descriptions use GPT-4o and Gemini API, which consumes about 800 dollars.
+  • Instructions with GUI-action chains take approximately 300 human-hour.
 A.2 Experiment details
-In our implementation, we utilize the internlm-7b variant of the SPHINX-X model, as detailed in [7].
-The pre-trained checkpoint for this model was sourced from the official repository mentioned in [21].
+In our implementation, we utilize the internlm-7b variant of the SPHINX model, as detailed in Liu et al.
+(2024).
+The pre-trained checkpoint for this model was sourced from the official repository mentioned in Team (2023).
 For image processing, the input images, each sized $1024\times 1024$, are segmented into sub-images.
-Visual features from these sub-images are extracted using two distinct visual encoders: DINOv2 [18] and ConvNext [25].
+Visual features from these sub-images are extracted using two distinct visual encoders: DINOv2 Oquab et al.
+(2024) and ConvNext Woo et al.
+(2023).
 To ensure compatibility in feature dimensions across different modules, linear projection layers are employed to align the channel dimensions.
-Regarding the model’s parameter settings, as outlined in Section 4, we configure the history window size to four.
+Regarding the model’s parameter settings, as outlined in Section 4.1, we configure the history window size to four.
 Additionally, we introduce a special token, <ICON>, specifically designed to identify interactive elements within the interface, strengthening the model’s interpretability and responsiveness to user interactions.
@@
 The fine-tuning was completed in four epochs.
+A.2.1 AndroidControl details
+AndroidControl splits the test set into four categories, i.e., in-domain data (IDD), category unseen, app unseen, and task unseen.
+Besides, the instructions are split into two levels.
+The high-level instructions are simply the goal of the task and the low-level instructions are the guides for each step during the task execution.
+Low-level tasks are easier for agents due to the fact that the agent only need to perform the action specified by the step-wise low-level instruction, while the high-level tasks require the agents to perform actions based on the overall instruction.
+For example, a step-wise low-level instruction is “click the profile element at the bottom right of the screen.”, compared with high-level instruction “Change my profile name to Unknown111”.
+A.2.2 AitW details
+Table 7: Human evaluation results on 5%-subsampled AitW test sub-set. “MMR” stands for “Mis-Match Random” subset.
+Subset Action Sources General Install G-Apps Single WebShopping Overall
+MMR SphAgent vs. Human 90.91 83.65 87.50 87.68 82.39 86.43
+AitW Anno. vs. Human 88.18 75.00 82.69 84.78 75.57 81.25
+Random SphAgent vs. Human 92.31 93.75 94.33 94.83 94.55 93.95
+AitW Anno. vs. Human 94.23 90.34 93.62 93.97 95.00 93.43
+We observe that there are several types of error cases in the AitW test set (see Figure 6).
+To assess the unreliability of the original AitW test set annotations, we hire human annotators to evaluate two subsets derived from the original test set.
+One subset is randomly chosen from episodes where SphAgent receives low accuracy (named Mis-Match Random), i.e., about 66% compared to the overall 78.72% shown in Table 5, indicating a high mismatch between inference results and AitW annotations.
+The other subset is randomly selected from the remaining episodes (named Random).
+Annotators first filter out repeated and redundant screenshots (see Figure 6(a)).
+For each remaining screenshot, the annotators then record the accuracies of SphAgent-inferred actions and the original AitW annotations versus our human evaluations.
+Figure 7 illustrates cases where annotators mark both the inferred action and the original annotation as correct, even though they interact with different elements.
+Table 7 presents the accuracy from human evaluation.
+The table indicates that the low SphAgent accuracy in the MMR subset are primarily due to unsatisfactory annotations (AitW Anno), which are used as ground truth during the evaluation.
+Both SphAgent and AitW Anno in the “MMR” subset have lower accuracies than those in the “Random” subset, highlighting that the tasks in the MMR subset are relatively more challenging.
+Furthermore, SphAgent achieves better human evaluation results than the original annotations, demonstrating its effectiveness and close alignment with human judgment.
+Comparing the results from Table 5 and Table 7, the notable differences in overall scores (i.e., 78.72%, 86.43%, and 93.95%) underscore the unreliability and misleading nature of the AitW test set annotations.
+(a) Repeating useless screenshots and actions.
+(b) Wrong task complete.
+Figure 6: Error cases in AitW test set. Red dots indicate the actions from the AitW annotations.
+Figure 7: Purple dots indicate the SphAgent inferred actions and red dots indicate the AitW annotations. Human annotators mark them both correct even though they are clicking different elements.
 A.3 More AMEX examples
-See Figure 9 for more examples of GUI interactive elements grounding and description.
-See Figure 10 and Figure 11 for more examples of instruction with GUI-action chains.
+See Figure 8 for more examples of GUI interactive elements grounding and description. See Figure 9 for more examples of instruction with GUI-action chains.
 A.4 Examples of Ethical Problems
@@
 Figure 9: More examples
-Figure 11: More examples
 Figure 10: Demonstration of anti-script mechanism where agents can enter the verification codes correctly.